OpenStudy (anonymous):

Write the explicit formula for the geometric sequence. a1 = -5 a2 = 20 a3 = -80?

OpenStudy (anonymous):

\(r=\frac{a_2}{a_1}=\frac{20}{-5}=-4\)

OpenStudy (anonymous):

can use \(a_n=-5\times (-4)^{n-1}\)